OpenStudy (anonymous):

Add: x-2/x-6 -4+ 4/x^2 a. -24+4x+22x^2-3x^3/x^2(x-6) b. 1/3+3 c. -4(x-2) d. -4(x-2)+4/(x-6) e. none of the above

OpenStudy (anonymous):

) Use disks: each disk has radius R = y = 1/x³, so V = ∫[a,b] πR² dx = π∫[1,3] dx/x^6 = (-π/5)x^-5 |[1,3] = (-π/5)(0.0041 - 1) ≈ 0.63 b) k affects the limits; you're asked to find where (-π/5)x^-5 |[1,k] = (-π/5)x^-5 |[k,3] -π/5 cancels; k^-5 - 1 = 3^-5 - k^-5 2k^-5 = 1 + 3^-5 = 1.0041 k ≈ 1.148 ← Check: 1.148^-5 - 1 = -0.498 3^-5 - 1.148^5 = -0.498 √ -0.498 * (-π/5) = 0.313 ≈ half the volume found in (a)

OpenStudy (anonymous):

Oh okay, hang on, I'll show you step by step.

OpenStudy (anonymous):

thank you @Aimee98 :)

OpenStudy (anonymous):

Write out the equation: \[x - \frac{ 2 }{ x } - 6 - 4 + \frac{ 4 }{ x^2 }\]

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

Then, simplify: \[x - \frac{ 2 }{ x } + \frac{ 4 }{ x^2 } - 10\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then lastly, cross multiply: \[x - \frac{ 2 (x - 2 + 5 x^2 }{ x^2 }\]

OpenStudy (anonymous):

So you answer?

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

one second finishing the math now.

OpenStudy (anonymous):

this is the same answer i got when i did the math. x^3-10x^2-2x+4/x^2

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

It looks excellent to me! When you got this last time was it wrong?